The surface of certain star has a surface temperature of about 5,050 K. The radius of the star is 7.2*108 m.

Firstly, calculate the total energy radiated by this star in each second expressed in watts. Note that the Stefan-Boltzmann constant is σ = 5.670*10-8 W/m2.

Secondly, take the log10 of the total calculated energy radiated in each second and input your answer into the answer box to one decimal.


1
Expert's answer
2021-10-07T08:48:19-0400
"P=4\pi R^2\sigma T^4\\
P=4\pi (7.2\cdot10^8)^2(5.67\cdot10^{-8})(5050)^4\\=2.4\cdot10^{26}\ W"